# Intrinsic Resolution of Organic Scintillators

### v1

@@ Abstract
Fluctuations in photon production within scintillators contribute to the overall energy resolution of scintillation detectors. While typically negligible, this contribution - known as intrinsic resolution (IR) - presents significant challenges for both experimental measurement and Monte Carlo simulation. Using the phenomenological model proposed in [1], we interpreted multiple datasets of IR measurements in organic scintillators. Although our results generally agree with the IR dataset examined in [1], a comprehensive evaluation of systematic uncertainties highlights the necessity for more detailed investigation. Based on our findings, we offer specific recommendations for future measurements.
@@ 1 Introduction
In our previous work [1], we proposed a simple single-parameter phenomenological model to describe the energy dependence of IR’s contribution to scintillation line width. The model assumes that the variance in photon production scales linearly with the mean photon yield, with its sole parameter representing the IR contribution at 1 MeV. Once calibrated using a setup with known photoelectron yield, this model allows for IR calculations across any detector configuration and energy within the model’s range of validity.
In this paper, we analyse energy resolution data for various organic scintillators reported by multiple research groups [4, 5, 6, 7, 8, 9, 10, 11, 12]. The primary objective of this analysis is to assess the consistency between experimental measurements and the theoretical model. Our results demonstrate that there are no statistically significant deviations from the model in the 10 keV to 4 MeV energy range. Although the model agrees well with the data, accurate parameter estimation is complicated by potential systematic uncertainties, particularly those relating to the calibration of the photomultiplier tubes (PMTs) used in the measurements.
@@ 2 Statistical description of the scintillations counter
In this section we review the classical description of the statistical properties of scintillation counter, following the formalism in [2], and introduces the framework for accounting for the contribution of the intrinsic resolution (IR). For simplicity, we restrict our analysis to point-like energy deposits.
@@
The average signal at the output of a three-stage cascade is [3]:
@@
For a non-ideal detector, the transfer probability $p$ may depend on multiple factors, including the photon’s point of origin, the details of the photon transfer process from that point to the photocathode, the angle and point of incidence on the photocathode and the efficiency of photoelectrons collection on the first dynode. Each set of possible transfer factors, "i", is characterised by a binomial distribution with a fixed average transfer probability $p_{i}$. The average transfer probability, $\overline{p}$, and its variance, $v(p)$, should be calculated across all possible sets of factors. Note that $v(p)$ is the variance of average probabilities $p_{i}$ here, rather than the variance of the random variable, $p$, as in the ideal case.
@@
Here $Q_{i}\equiv\frac{p_{i}}{\overline{p}}\overline{Q}$ and $w_{i}$ is the weight of the corresponding set of transfer factors. Noting that $\sum_{i}w_{i}Q_{i}^{2}=\overline{Q}(1+v(p))$ and $\sum_{i}w_{i}Q_{i}=\overline{Q}$, we can rewrite (5) as
@@
$v(Q)\equiv\frac{Var(Q)}{\overline{Q}^{2}}=\frac{1+v_{1}}{Q}+v(p)+\\ +(1+v(p))\left[v(N_{\text{ph}})-\frac{1}{\overline{N_{\text{ph}}}}\right].$ (7)
@@
As can be seen from equation (8), the additional variance due to the IR contributes in the same way as the statistical contribution, $\frac{1+v_{1}}{Q}$, as they have the same dependence on $Q$. Therefore, within the framework of the considered formalism, the IR effect imitates the statistical smearing of the charge signal. This effect is more pronounced for detectors with a high photoelectron yield. At an energy deposit of 1 MeV (i.e. at $Q=Q_{1}$), the IR contribution should be compared to the value $\frac{1+v_{1}^{\text{Det}}}{Q_{1}}$: the higher $Q_{1}$, the higher the relative contribution of the IR.
@@
Taking into account the low values of $v_{1}^{\text{int}}$, which are typically in the region of ×$\times 10^{-4}$ units (see [1] and table 3), it is clear that an experiment involving a high yield of photoelectrons would be more suitable for measuring $v_{1}^{\text{int}}$. For example, with an expected IR of 2% at 1 MeV ($v_{1}^{\text{int}}=4\times 10^{-4}$) and a relative yield of 100 p.e./MeV, the IR contribution is $0.0410^{2}\times 0.02^{2}=$ 0.04, which is small compared to $1+v_{1}$ for any $v_{1}$. For a relative yield of 500 p.e./MeV and the same $v_{1}^{\text{int}}$, the IR contribution reaches 0.2, making it easier to measure. Another important factor in this type of measurement is the PMT resolution in a single p.e. mode, characterised by the parameter $v_{1}$. As it contributes in the same way as the IR, any uncertainty in its definition translates into the uncertainty in the IR effect. Due to the complex structure of the PMT response and the presence of noise, measuring the $v_{1}$ itself is a complicated task.
@@
In this paper, we will use the notation $Q$ for the measured charge, while keeping the calibration parameter "c".
We would like to highlight the implicit assumption of linearity in the conversion from $N_{\text{ph}}$ in $Q$, and the explicit one concerning the monoenergetic nature of the source. If the source is not monoenergetic, the term of the type $\frac{\Delta Q^{2}}{Q^{2}}$ should be taken into account, where $\Delta{Q}^{2}$ corresponds to the source energy variance.
@@
where $LY$ is the relative p.e. yield at 1 MeV, the electron energy, $E$, is expressed in MeV, and the function $f_{\text{NL}}(E)$ accounts for possible nonlinearities in the conversion of energy into light. Since all the measurements considered in this paper are reported on an energy scale, we use the approximation $v(Q)\simeq v(E)$. This approximation is justified because the authors apparently used the inverse approximation, $v(E)\simeq v(Q)$, implicitly, to obtain the energy resolution.
@@ 3 The intrinsic resolution for gamma’s
When calculating the sum $\sum_{i}Q_{e}(i)^{2}$, we used the fact of statistical independence of the individual contributions. In other words, the contribution of transfer non-uniformity can be applied to the total light collected if the detector is uniform, i.e. if all the light produced by Compton electrons exhibits the same transfer properties.
@@ 4 Absolute PMT calibration and other sources of systematics in IR measurements
We identified a systematic issue in PMT calibration, affecting nearly all measurements reporting the relative light yield (photoelectrons per MeV). Specifically, in [4][6], the peak of the SER distribution is incorrectly associated with one photoelectron, rather than the true mean value of the SER  [^2] We prefer to use the term ”main peak” to describe the feature. Additionally, we use ”(main) peak position” instead of ”most probable value” because, for spectra with a strong underamplified branch, the most probable value may approach zero.. Furthermore, in some of these studies, the relative variance $v_{1}$ of the SER is either calculated from the peak of the SER distribution, ignoring deviations from a Gaussian [6], or arbitrarily assumed to be $v_{1}=0.1$ without presenting any measurements [4]. Additionally, the same research group in [20] attributes discrepancies between single-electron and multi-electron PMT calibrations as a "new source of error" in scintillation measurements. However, the actual cause is the substituting of the true mean and full variance of the SER with values derived solely from the "normal multiplication" peak in low-gain measurements.
Taking the issue of calibration into account, we attempted to estimate the associated systematics for each set of measurements. The relevant information concerning the setup of the measurements considered in this paper is summarised in Table 1. All measurements except from those of Borexino and [7] are performed using cylindrical samples (cylindrical quartz vials in the case of LS).
@@ 5 Measurements with BC408 and EJ301 commercial scintillators
In this section we analyse the data from the publication [4], and all the information below originates from this article unless clearly stated otherwise. Our main interest lies in the results of the measurement of the intrinsic resolution of commercial EJ301 liquid scintillator and BC408 plastic scintillator. The setup of the experiment was intended for the intrinsic resolution measurement. It provides an excellent light yield, which is the most important quantity in this kind of measurements, as discussed above.
@@
We subtracted the uncertainty associated with $LY$ from the data uncertainties.
To check the compatibility of the experimental data with the model (10) we fitted the data to a function derived from (9):
$v_{\text{int}}(E)=\frac{1+v_{1}}{LY\cdot E\cdot f_{\text{NL}}(E)}+\\ +v(p)+(1+v(p))\frac{f_{\text{NL}}(1)}{f_{\text{NL}}(E)\cdot E}v_{1}^{\text{int}}.$ (16)
As the IR data is presented as a function of energy, we substituted the ratio $\frac{Q_{1}}{Q}$ by its equivalent $\frac{f_{\text{NL}}(1~{}\text{MeV})}{f_{\text{NL}}(E)}$. Here $f_{\text{NL}}$ is the nonlinearity function from Fig.6 of the paper [4]. The values in Fig.6 [4] are normalised to unity at 447 keV. To use the data in our analysis, we fitted it with a smooth curve and extracted the value for $f_{\text{NL}}(1~{}\text{MeV})$. All other data can be used in their original form. The uncertainty of the $f_{\text{NL}}(E)$ function was neglected in the analysis. In the intrinsic resolution extraction procedure, the term $v(p)$ ($\delta_{p}$ in the authors’ notation) was omitted. As stated in the article: "The … term is linked to the light collection at the photocathode and the transfer of photoelectrons to the first dynode. It is assumed that these effects do not introduce substantial spread to the measured pulse width in modern PMTs." While the term is indeed small, the data on the right-hand side of the energy range clearly show a constant "supporting" term. This is why we left the term free in the analysis.
A straightforward fit of the data to the function (16) yields a very small value of $\chi^{2}$, with a reduced value $\chi^{2}/n.d.f.\simeq 0.3$. This indicates that the errors are overestimated. In fact, the original uncertainties in Fig.7 [4] take two values: 10% at lower energies and 7.5% at higher energies. We suggest that this is a generic type of uncertainty assigned to the instrument. To obtain more realistic probabilistic estimates of the results, we scaled the resulting uncertainties by a factor of 0.5, yielding reduced $\chi^{2}$ values close to 1 for this set of measurements.
@@
Figure 1: The energy resolution of the EJ301 fitted using equation (16). The LY and $v_{1}$ parameters are fixed at the values provided in the paper [4]. The figure shows the contributions of statistical fluctuations, IR and the (flat) contribution of light transfer non-uniformity.
@@
Figure 2: The energy resolution of the BC408 fitted using (16). The LY and $v_{1}$ parameters are fixed at the values provided in the paper.
@@
The possible bias in PMT calibration (i.e. deviation of the factor "c" from unity) is not discussed at all in the article [4]. The uncertainty of the $Q$ measurement can be derived from the light yield estimate. The reported LY for the EJ301 scintillator is $2850\pm 150$ p.e./MeV (5.4%) and $2770\pm 150$ (5.3%) for the BC408 one. As discussed above, the single p.e. response was characterised using the position of the peak in the single p.e. in [4], thus introducing a bias in the measurement (i.e. $c\neq 1$). The peak position does not necessarily coincide with the true single p.e. average value. In the presence of a strong fraction of the underamplified  [^3] In a book [15] these signals are called undersized, we prefer the term ”underamplified” signals, the difference could be significant. The same applies to the evaluation of the $v_{1}$ value from fitting a Gaussian to the peak. As will be shown below on an example of another measurement, the difference could be quite significant. The value of $v_{1}=0.1$ is mentioned for the case of the XP5212B PMT used in measurements with the EJ301 liquid scintillator. Apparently, the same value was used for the BC408 scintillator measurements with a different PMT, although this is not explicitly stated in the paper.
@@ 6 Measurements with the EJ200 plastic scintillator
where $E$ is energy measured in MeV. For our purposes, we renormalised the function to unity at 1 MeV by dividing $f_{\text{NL}}(E)$ by $f_{\text{NL}}(1)=1.02$. As the non-linearity was studied using a set of data with minimum energy of $E\simeq$ 100 keV, doubts remain whether (18) can be extrapolated to lower energies, given that at least four data points in the resolution plot are below 100 keV. Thus, we applied both the interpolation (18) and the results of the BC408 non-linearity measurements (an analogue of EJ200) from [4]? and found no differences.
@@
A straightforward fit using the expression (16) apparently underestimates the values at low energies. This is due to the energy binning used in the analysis ($\Delta E$=10 keV) at the germanium detector side. To account for the width of the energy bin, we added the term $\frac{{\Delta E}^{2}}{12\cdot(f_{\text{NL}}\cdot E)^{2}}$ with a fixed $\Delta E$ to the expression (16), assuming a rectangular distribution of energy within the 10 keV energy bin. The light yield of 2500 p.e./MeV is quoted in the caption of Fig.7 [8], though no details of the measurements are presented in the text. The "statistical term" in Fig.7 [8] is plotted without considering the PMT response width $v_{1}$. The setup (a 50 mm diameter plastic scintillator with XP5500 PMT) is similar to the measurements with the BC408 considered above. Taking into account that the manufacturers state the light output of 64% anthracene for both scintillators, we can reasonably assume the same LY in both measurements ($2770\pm 150$ p.e./MeV).
Figure 3: The single electron response of PMT XP5500 from [17].
An example of the single-electron response (SER) of the Photonis XP5500 photomultiplier tube (from [17], Fig. 7, the one taken at +35 C) is reproduced in Fig. 3. The SER spectrum exhibits a pronounced underamplified signal branch and a long tail of high-amplitude events, the latter are likely to be a contribution from multi-electrons response. No further experimental details or numerical estimates of SER parameters are provided in [17]. To quantify the observed distribution, we calculated the mean charge $q_{1}=61.1$ and the relative variance $v_{1}=0.42$ for the full spectrum. The main peak was also fitted with a Gaussian, yielding $q_{1}^{\text{peak}}=63.8$ and $v_{1}^{\text{peak}}=0.20$. The minimal discrepancy between $q_{1}$ and $q_{1}^{\text{peak}}$ in the presence of a strong underamplified component suggests that the measured distribution does not represent a "pure" SER, but rather a response to a low-intensity light source with pedestal subtracted. To obtain a more realistic estimate of the parameters, we calculated $q_{1}=50.0$ from the distribution cut at $\simeq 2\cdot q_{1}=100$. The corresponding $c$ value is $c=0.78$ and the light yield is then $LY=3550$, noticeably high compared to $2770\pm 150$ p.e./MeV. Nevertheless, the new value is close to the value of $LY=4050\pm 200$ reported for a very similar setup with the same type of PMT [18]. The relative variance of the truncated distribution is $v_{1}=0.31$
We also reconstructed a reasonable value for the relative variance, $v_{1}$, differently. First, we assumed that the spectrum consists of a peaked feature with a probability $p$, and that the remaining $(1-p)$ fraction do not contribute significantly to the signal, producing values with an average charge $q_{\text{U}}$ close to 0. The assumption is reasonable, as can be seen from Fig. 3. The true average value of the single electron spectrum is then:
@@
We set $v_{1}=0.33$ and increased the LY by 22%, the approximate difference between the peak and the average value. We also assumed 10% uncertainty for LY. The results of the fit are presented in Fig. 4.
Figure 4: The energy resolution of the EJ200 fitted with (16). The choice of the LY and $v_{1}$ parameters is explained in the text. The figure shows the contributions of statistical fluctuations, IR and (flat) contribution of light transfer non-uniformity.

### v2

@@ Abstract
Fluctuations in photon production within scintillators contribute to the overall energy resolution of scintillation detectors. While typically negligible, this contribution - known as intrinsic resolution (IR) - presents significant challenges for both experimental measurement and Monte Carlo simulation. Using the phenomenological model proposed in [1], we interpreted multiple datasets of IR measurements in organic scintillators. Although my results generally agree with the IR dataset examined in [1], a comprehensive evaluation of systematic uncertainties highlights the necessity for more detailed investigation. Based on my findings, we offer specific recommendations for future measurements.
@@ 1 Introduction
In my previous work [1], we proposed a simple single-parameter phenomenological model to describe the energy dependence of IR’s contribution to scintillation line width. The model assumes that the variance in photon production scales linearly with the mean photon yield, with its sole parameter representing the IR contribution at 1 MeV. Once calibrated using a setup with known photoelectron yield, this model allows for IR calculations across any detector configuration and energy within the model’s range of validity.
In this paper, we analyse energy resolution data for various organic scintillators reported by multiple research groups [3, 4, 5, 6, 7, 8, 9, 10, 11]. The primary objective of this analysis is to assess the consistency between experimental measurements and the theoretical model. My results demonstrate that there are no statistically significant deviations from the model in the 10 keV to 4 MeV energy range. Although the model agrees well with the data, accurate parameter estimation is complicated by potential systematic uncertainties, particularly those relating to the calibration of the photomultiplier tubes (PMTs) used in the measurements.
@@ 2 Statistical description of the scintillations counter
In this section we review the classical description of the statistical properties of scintillation counter, following the formalism in [12], and introduces the framework for accounting for the contribution of the intrinsic resolution (IR). For simplicity, we restrict the analysis to point-like energy deposits.
@@
In the general case, the average signal at the output of a three-stage cascade is [13]:
@@
For a non-ideal detector, the transfer probability $p$ may depend on multiple factors, including the photon’s point of origin, the details of the photon transfer process from that point to the photocathode, the angle and point of incidence on the photocathode and the efficiency of photoelectrons collection on the first dynode. Each set of possible transfer factors, "i", is characterised by a binomial distribution with a fixed average transfer probability $p_{i}$. The average transfer probability, $\overline{p}$, and its variance, $v(p)$, should be calculated across all possible sets of factors. Note that notation $v(p)$ refers to the variance of the values of average probabilities $p_{i}$, in contrast to the variance of the random binomial variable with a fixed average probability $p$ in the previously considered case.
@@
Here $Q_{i}\equiv\frac{p_{i}}{\overline{p}}\overline{Q}$ and $w_{i}$ is the weight of the corresponding set of transfer factors. Equation (4) was used to develop $\overline{Q_{i}^{2}}$. Noting that $\sum_{i}w_{i}Q_{i}^{2}=\overline{Q}^{2}(1+v(p))$ and $\sum_{i}w_{i}Q_{i}=\overline{Q}$, we can rewrite (5) as
@@
$v(Q)\equiv\frac{Var(Q)}{\overline{Q}^{2}}=\frac{1+v_{1}}{\overline{Q}}+v(p)+\\ +(1+v(p))\left[v(N_{\text{ph}})-\frac{1}{\overline{N_{\text{ph}}}}\right].$ (7)
@@
As can be seen from equation (8), the additional variance due to the IR contributes in the same way as the statistical contribution, $\frac{1+v_{1}}{Q}$, as they have the same dependence on $Q$. Therefore, within the framework of the considered formalism, the IR effect imitates the statistical smearing of the charge signal. This effect is more pronounced for detectors with a high photoelectron yield. At an energy deposit of 1 MeV (i.e. at $Q=Q_{1}$), the IR contribution $v_{1}^{\text{int}}$ should be compared to the value $\frac{1+v_{1}^{\text{Det}}}{Q_{1}}$: the higher $Q_{1}$, the higher the relative contribution of the IR.
@@
Taking into account the low values of $v_{1}^{\text{int}}$, which are typically of the order of $10^{-4}$ (see [1] and table 3), it is clear that an experiment involving a high yield of photoelectrons would be more suitable for measuring $v_{1}^{\text{int}}$. For example, with an expected IR of 2% at 1 MeV ($v_{1}^{\text{int}}=4\times 10^{-4}$) and a relative yield of 100 p.e./MeV, the IR contribution is $10^{2}\times 0.02^{2}=0.04$, which is small compared to $1+v_{1}$ for any $v_{1}$. For a relative yield of 500 p.e./MeV and the same $v_{1}^{\text{int}}$, the IR contribution reaches 0.2, making it easier to measure. Another important factor in this type of measurement is the PMT resolution in a single p.e. mode, characterised by the parameter $v_{1}$. As it contributes in the same way as the IR, any uncertainty in its definition translates into the uncertainty in the IR effect. Due to the complex structure of the PMT response and the presence of noise, measuring the $v_{1}$ itself is a complicated task.
@@
In this paper, I will use the notation $Q$ for the measured charge, while keeping the calibration parameter "c".
I would like to highlight the implicit assumption of linearity in the conversion from $N_{\text{ph}}$ in $Q$, and the explicit one concerning the monoenergetic nature of the source. If the source is not monoenergetic, the term of the type $\frac{\Delta Q^{2}}{Q^{2}}$ should be taken into account, where $\Delta{Q}^{2}$ corresponds to the source energy variance.
@@
where $LY$ is the relative p.e. yield at 1 MeV, the electron energy, $E$, is expressed in MeV, and the function $f_{\text{NL}}(E)$ accounts for possible nonlinearities in the conversion of electron kinetic energy into light. Note that similar relations can be used for other particles ($\gamma$, protons, $\alpha$-particles, etc.) with their corresponding light yield ($LY$) and $f_{\text{NL}}(E)$ functions.
Since all the measurements considered in this paper are reported on an energy scale, we use the approximation $v(Q)\simeq v(E)$. This approximation is justified because the authors apparently used the inverse approximation, $v(E)\simeq v(Q)$, implicitly, to obtain the energy resolution.
@@ 3 The intrinsic resolution for gamma’s
When calculating $\sum_{i}Q_{e}(i)^{2}$, we used the fact of statistical independence of the individual contributions. In other words, the contribution of transfer non-uniformity can be applied to the total light collected if the detector is uniform, i.e. if all the light produced by Compton electrons exhibits the same transfer properties.
@@ 4 Absolute PMT calibration and other sources of systematics in IR measurements
I identified a systematic issue in PMT calibration, affecting nearly all measurements reporting the relative light yield (photoelectrons per MeV). Specifically, in [3][5], the peak of the SER distribution is associated with one photoelectron, rather than the true mean value of the SER  [^2] I prefer to use the term ”main peak” to describe the feature. Additionally, I use ”(main) peak position” instead of ”most probable value” because, for spectra with a strong underamplified branch with exponential behaviour, the most probable value may approach zero.. Furthermore, in some of these studies, the relative variance $v_{1}$ of the SER is either calculated from the peak of the SER distribution, ignoring deviations from a Gaussian [5], or arbitrarily assumed to be $v_{1}=0.1$ without presenting any measurements [3]. Additionally, the same research group in [17] attributes discrepancies between single-electron and multi-electron PMT calibrations as a "new source of error" in scintillation measurements. However, the actual cause is the substituting of the true mean and full variance of the SER with values derived solely from the "normal multiplication" peak in low-gain measurements.
Taking the issue of calibration into account, I attempted to estimate the associated systematics for each set of measurements. The relevant information concerning the setup of the measurements considered in this paper is summarised in Table 1. All measurements except from those of Borexino and [6] are performed using cylindrical samples (cylindrical quartz vials in the case of LS).
@@ 5 Measurements with BC408 and EJ301 commercial scintillators
In this section I analyse the data from the publication [3], and all the information below originates from this article unless clearly stated otherwise. Our main interest lies in the results of the measurement of the intrinsic resolution of commercial EJ301 liquid scintillator and BC408 plastic scintillator. The setup of the experiment was intended for the intrinsic resolution measurement. It provides an excellent light yield, which is the most important quantity in this kind of measurements, as discussed above.
@@
I subtracted the uncertainty associated with $LY$ from the data uncertainties.
To check the compatibility of the experimental data with the model (10) I fitted the data to a function derived from (9):
$v_{\text{int}}(E)=\frac{1+v_{1}}{LY\cdot E\cdot f_{\text{NL}}(E)}+\\ +v(p)+(1+v(p))\frac{f_{\text{NL}}(1~\text{MeV})}{f_{\text{NL}}(E)\cdot E}v_{1}^{\text{int}}.$ (16)
As the IR data is presented as a function of energy, I substituted the ratio $\frac{Q_{1}}{Q}$ by its equivalent $\frac{f_{\text{NL}}(1~\text{MeV})}{f_{\text{NL}}(E)}$. Here $f_{\text{NL}}$ is the nonlinearity function from Fig.6 of the paper [3]. The values in Fig.6 [3] are normalised to unity at 447 keV. To use the data in the analysis, I fitted it with a smooth curve and extracted the value for $f_{\text{NL}}(1~\text{MeV})$. All other data can be used in their original form. The uncertainty of the $f_{\text{NL}}(E)$ function was neglected in the analysis. In the intrinsic resolution extraction procedure, the term $v(p)$ ($\delta_{p}$ in the authors’ notation) was omitted. As stated in the article: "The … term is linked to the light collection at the photocathode and the transfer of photoelectrons to the first dynode. It is assumed that these effects do not introduce substantial spread to the measured pulse width in modern PMTs." While the term is indeed small, the data on the right-hand side of the energy range clearly show a constant "supporting" term. This is why we left the term free in the analysis.
A straightforward fit of the data to the function (16) yields a very small value of $\chi^{2}$, with a reduced value $\chi^{2}/n.d.f.\simeq 0.3$. This indicates that the errors are overestimated. In fact, the original uncertainties in Fig.7 [3] take two values: 10% at lower energies and 7.5% at higher energies. I suggest that this is a generic type of uncertainty assigned to the instrument. To obtain more realistic probabilistic estimates of the results, I scaled the resulting uncertainties by a factor of 0.5, yielding reduced $\chi^{2}$ values close to 1 for this set of measurements.
@@
Figure 1: The energy resolution of the EJ301 fitted using equation (16). The source of the data is Fig.7 of [3], the LY and $v_{1}$ parameters are fixed at the values provided in the paper [3]. The figure shows the contributions of statistical fluctuations, IR and the (flat) contribution of light transfer non-uniformity.
@@
Figure 2: The energy resolution of the BC408 fitted using (16). The source of the data is Fig.4 of [3]), the LY and $v_{1}$ parameters are fixed at the values provided in the paper.
@@
The possible bias in PMT calibration (i.e. deviation of the factor "c" from unity) is not discussed at all in the article [3]. The uncertainty of the $Q$ measurement can be derived from the light yield estimate. The reported LY for the EJ301 scintillator is $2850\pm 150$ p.e./MeV (5.4%) and $2770\pm 150$ (5.3%) for the BC408 one. As discussed above, the single p.e. response was characterised using the position of the peak in the single p.e. in [3], thus introducing a bias in the measurement (i.e. $c\neq 1$). The peak position does not necessarily coincide with the true single p.e. average value. In the presence of a strong fraction of the underamplified  [^3] In a book [16] these signals are called undersized, I prefer the term ”underamplified” signals, the difference could be significant. The same applies to the evaluation of the $v_{1}$ value from fitting a Gaussian to the peak. As will be shown below on an example of another measurement, the difference could be quite significant. The value of $v_{1}=0.1$ is mentioned for the case of the XP5212B PMT used in measurements with the EJ301 liquid scintillator. Apparently, the same value was used for the BC408 scintillator measurements with a different PMT, although this is not explicitly stated in the paper.
@@ 6 Measurements with the EJ200 plastic scintillator
where $E$ is energy measured in MeV. For our purposes, we renormalised the function to unity at 1 MeV by dividing $f_{\text{NL}}(E)$ by $f_{\text{NL}}(1)=1.02$. As the non-linearity was studied using a set of data with minimum energy of $E\simeq$ 100 keV, doubts remain whether (18) can be extrapolated to lower energies, given that at least four data points in the resolution plot are below 100 keV. Thus, we applied both the interpolation (18) and the results of the BC408 non-linearity measurements (an analogue of EJ200) from [3] and found no differences.
@@
A straightforward fit using the expression (16) apparently underestimates the values at low energies. This is due to the energy binning used in the analysis ($\Delta E$=10 keV) at the germanium detector side. To account for the width of the energy bin, we added the term $\frac{{\Delta E}^{2}}{12\cdot(f_{\text{NL}}\cdot E)^{2}}$ with a fixed $\Delta E$ to the expression (16), assuming a rectangular distribution of energy within the 10 keV energy bin. The light yield of the setup is not discussed in the text of [7]. The value of 2500 p.e./MeV is used in the caption of Fig.7 [7] for illustration purpose, but the "statistical term" in Fig.7 [7] is plotted without considering the PMT response width $v_{1}$. Because the setup used (a 50 mm diameter plastic scintillator with XP5500 PMT) is similar to the measurements with the BC408 considered above and taking into account that the manufacturers state the light output of 64% anthracene for both scintillators, we can reasonably assume the same LY in both measurements, i.e. $2770\pm 150$ p.e./MeV.
Figure 3: The single electron response of PMT XP5500 from [18] (Fig. 7, the one taken at +35 C).
An example of the single-electron response (SER) of the Photonis XP5500 photomultiplier tube is reproduced in Fig. 3. The SER spectrum exhibits a pronounced underamplified signal branch and a long tail of high-amplitude events, the latter are likely to be a contribution from multi-electrons response. No further experimental details or numerical estimates of SER parameters are provided in [18]. To quantify the observed distribution, we calculated the mean charge $q_{1}=61.1$ and the relative variance $v_{1}=0.42$ for the full spectrum. The main peak was also fitted with a Gaussian, yielding $q_{1}^{\text{peak}}=63.8$ and $v_{1}^{\text{peak}}=0.20$. The minimal discrepancy between $q_{1}$ and $q_{1}^{\text{peak}}$ in the presence of a strong underamplified component suggests that the measured distribution does not represent a "pure" SER, but rather a response to a low-intensity light source with pedestal subtracted. To obtain a more realistic estimate of the parameters, we calculated $q_{1}=50.0$ from the distribution cut above 100, corresponding to $\simeq 2\cdot q_{1}$. The corresponding $c$ value is then $c=0.78$ (50/63.8) and the light yield is $LY=3550$, noticeably high compared $2770\pm 150$ p.e./MeV. Nevertheless, the new value is more close to the value of $LY=4050\pm 200$ reported for a very similar setup with the same type of PMT [19]. The relative variance of the truncated distribution is $v_{1}=0.31$
I also reconstructed a value for the relative variance, $v_{1}$, differently. First, I assumed that the spectrum consists of a peaked feature with a probability $p$, and that the remaining (1p)(1-p) fraction do not contribute significantly to the signal, producing values with an average charge $q_{\text{U}}$ close to 0. The assumption is reasonable, as can be seen from Fig. 3. The true average value of the single electron spectrum is then:
@@
I set $v_{1}=0.33$ and increased the LY by 22%, the approximate difference between the peak and the average value. I also accepted 10% uncertainty for LY. The results of the fit are presented in Fig. 4.
Figure 4: The energy resolution of the EJ200 fitted with (16). The source of data is Fig.7 of [7]. The choice of the LY and $v_{1}$ parameters is explained in the text. The figure shows the contributions of statistical fluctuations, IR and (flat) contribution of light transfer non-uniformity.